A plane travels 1200 mi with the wind in 2 hours. The return trip against the wind takes 3 hours. Find the rate of speed of the plane in still air.
*Only trying to find rate of speed of plane in still air, not wind speed*

Answer by TimothyLamb(4379)

s = d / t
s + w = 1200 / 2 = 600
s - w = 1200 / 3 = 400
---
s = 600 - w
s = 400 + w
600 - w = 400 + w
200 = 2w
w = 100
Answer:
s = 500
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php